OpenStudy (anonymous):

solve w^4-4w²-2=0 cant figure out what i am doing wrong u=w² u²-4u-2=0 u=2-(sqrt)6 or u=2+(sqrt)6 w²=2-(sqrt)6 or w²=2+(sqrt)6 w=(sqrt)2+(sqrt)6, -(sqrt)2+(sqrt)6 tell me what i am doing wrong and how to fix it plz

OpenStudy (anonymous):

nothing wrong at the beginning. you get \[w^2=2+\sqrt{6}\] \[w=\pm \sqrt{2+\sqrt{6}}\]

OpenStudy (anonymous):

it is your final calculation that is off. you only need to take the square root of your solution for \[w^2\]

OpenStudy (anonymous):

now \[2-\sqrt{6}<0\] so if you are working only with real numbers this answer gives no solution for w

OpenStudy (anonymous):

did this make sense? the only mistake you made was taking the square root of \[2+\sqrt{6}\]

OpenStudy (anonymous):

shes an idiot

OpenStudy (anonymous):

thats whats off

OpenStudy (anonymous):

you have to put the square root sign over the whole expression

OpenStudy (anonymous):

w=sqrt [ 2+sqrt [6] ]